OpenStudy (anonymous):

Use the three steps to solve the problem. Express the answers using decimals. It took John 12 hours riding his bike to make the round trip to his uncle's. If he averaged 20 mph out and 30 mph back, how long did he travel each way? (Round answer to nearest tenth.)

OpenStudy (anonymous):

Refer to the attachment developed with Mathematica v9.

OpenStudy (mokeira):

First find the distance traveled \[12=\frac{ d }{20 } + \frac{ d }{ 30 }\] multiply everything by 60 to remove denominators. \[60( 12=\frac{ d }{20 }+\frac{ d }{ 30 })\] \[720=3d+20\] \[d=144m\] Now you can find time for each Time to: \[144m \div 20mph\] \[=7.2h\] Time from: \[144m \div 30mph\] \[=4.8h\]

OpenStudy (anonymous):

Refer to the solution which includes the distance traveled in each direction.